OpenStudy (somy):

Why is it that i always can't see full drawings posted?? i mean i can only see half of it >_< can anybody tell me why is that?

OpenStudy (somy):

like this for example

OpenStudy (somy):

@e.mccormick @thomaster

thomaster (thomaster):

It gets cut of on the right side, always. It's a known bug. Just don't get too close to the right border when you use the draw tool :)

OpenStudy (anonymous):

lol this is wht i wrote \[\sin \alpha = \frac{h}{s}\] we need to find h ( for gpe) \[h= \sin \alpha \times s\]

OpenStudy (somy):

ah lol thanx no worries i just used that pic to show what usually happens :) @thushananth01 and thank you for info @thomaster :D